OpenStudy (anonymous):

If the outliers are not included, what is the mean of the data set? 72, 82, 3, 55, 62, 73, 90, 80, 151, 59, 46, 31 (2 points) 59 65 67 73

OpenStudy (anonymous):

i think the answer is b or a

TheSmartOne (thesmartone):

Do you know what mean is?

TheSmartOne (thesmartone):

\(\bf\huge~~~~~\color{#ff0000}{W}\color{#ff2000}{e}\color{#ff4000}{l}\color{#ff5f00}{c}\color{#ff7f00}{o}\color{#ffaa00}{m}\color{#ffd400}{e}~\color{#bfff00}{t}\color{#80ff00}{o}~\color{#00ff00}{O}\color{#00ff40}{p}\color{#00ff80}{e}\color{#00ffbf}{n}\color{#00ffff}{S}\color{#00aaff}{t}\color{#0055ff}{u}\color{#0000ff}{d}\color{#2300ff}{y}\color{#4600ff}{!}\color{#6800ff}{!}\color{#8b00ff}{!}\\\bf ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~Made~by~TheSmartOne\) Hey there!!! Since you are new here, read this legendary tutorial for new OpenStudiers!! http://openstudy.com/study#/updates/543de42fe4b0b3c6e146b5e8 @helpmeimagymnast

OpenStudy (anonymous):

yeah

TheSmartOne (thesmartone):

So did you add up all the numbers and divide by how mane there are?

OpenStudy (anonymous):

you add up all of the numbers and then divide by the amount of numbers

jagr2713 (jagr2713):

great job eliminating @helpmeimagymnast

TheSmartOne (thesmartone):

Good Job :)

jagr2713 (jagr2713):

\[\huge{\color{purple}{\textbf{W}} \color{orange}{\cal{E}} \color{green}{\mathbb{L}} \color{blue}{\mathsf{C}} \color{maroon}{\rm{O}} \color{red}{\tt{M}} \color{gold}{\tt{E}} \space \color{orchid}{\mathbf{T}} \color{Navy}{\mathsf{O}} \space \color{OrangeRed}{\boldsymbol{O}} \color{Olive}{\mathbf{P}} \color{Lime}{\textbf{E}} \color{DarkOrchid}{\mathsf{N}} \color{Tan}{\mathtt{S}} \color{magenta}{\mathbb{T}} \color{goldenrod}{\mathsf{U}} \color{ForestGreen}{\textbf{D}} \color{Salmon}{\mathsf{Y}} \ddot \smile }\]

OpenStudy (anonymous):

i know that 151 is an outliner but im not sure if 3 is

jagr2713 (jagr2713):

pick one a or b and we will go from there :D

OpenStudy (anonymous):

i say b

jagr2713 (jagr2713):

CORRECT

TheSmartOne (thesmartone):

Correct! :)

OpenStudy (anonymous):

only if 3 is an outliner.....really im confused can you explain it

jagr2713 (jagr2713):

so first we put the values in ascending order.

jagr2713 (jagr2713):

3 31 46 55 59 62 72 73 80 82 90 151

OpenStudy (anonymous):

ok

jagr2713 (jagr2713):

wait u got this @TheSmartOne

OpenStudy (anonymous):

then i find the median, Q1, and Q3

TheSmartOne (thesmartone):

outliner are the numbers that are the smallest and the largest in the qroup of numbers

jagr2713 (jagr2713):

yea we find the median :D

OpenStudy (anonymous):

then i minus Q3 from Q1 which I got 81-49

jagr2713 (jagr2713):

81 is correct but 49 isnt D:, which is Q3 and Q1

OpenStudy (anonymous):

so then what is Q1?

jagr2713 (jagr2713):

you were close though

OpenStudy (anonymous):

what is Q1? im not sure now because I though it was 49

jagr2713 (jagr2713):

Find the first quartile Q1: The number of values below the median is: 3 31 46 55 59 62 Find the median of these values n=6 (n+1)/2 = 3.5 3.5 lies between 3 and 4 Median of the 6 values = (46+55)/2 = 50.5 Q1 = 50.5

jagr2713 (jagr2713):

you get it?

OpenStudy (anonymous):

oooooo i get it now I had the wrong number... haha thank you for you time @jagr2713 and @TheSmartOne

jagr2713 (jagr2713):

anytime buddy :D

OpenStudy (anonymous):

could you check one for me?

OpenStudy (anonymous):

another i mean?

jagr2713 (jagr2713):

ummm ill try :D

OpenStudy (anonymous):

Given the box plot, will the mean or the median provide a better description of the center? (1 point)

OpenStudy (anonymous):

here ill open another question box @jagr2713

jagr2713 (jagr2713):

ok

OpenStudy (anonymous):

hang on im having trouble

jagr2713 (jagr2713):

take your time :D